# First investigation of void statistics in numerical relativity simulations

### v1

@@ 4.5 Boundary distance profiles
We also produce profiles using this approach, using the set of centres of wall cells adjacent to the interior cells of a given void as the boundary for that void, and define distance from the boundary as the distance to the closest point in that set.
Figs. 11 and 12 show the averaged profile of $\delta$ top panels), $\Theta$ (middle panels), and $\tensor\tensor[]{\mathcal{R}}{}$ (bottom panels) in the two simulations as a function of distance from the void boundary. The extremum that appears in the walls for each variable has greater magnitude than for the corresponding stacked spherical profile. Likewise, the gradient near the walls is greater. These properties are to be expected, as the walls are all aligned at distance $d=0$, rather than scattered at varying radial distances around $r=R_{\mathrm{eff}}$.
@@ 5 Conclusions
In this work, we explored void statistics in non-linear general relativity for the first time. Profiles exhibiting how expansion and curvature fluctuate inside “empty” regions demonstrate markedly different results to common assumptions in Newtonian simulations.
We have implemented the Watershed Einstein Topography (WET) void finder for use with simulation data output by the Einstein Toolkit. Our simulations approximate matter as a continuous fluid, which necessitates differences in the implementation of this void finder compared to publicly-available void finders.
Having identified the voids in two simulations of different physical resolution, we studied the void size function, void fraction, and stacked radial density profiles. We fit HSW profiles (Hamaus et al., 2014) to our stacked radial density profiles for each simulation. As discussed in Section 4.4, we alter this process by fitting a parameter $R_{\mathrm{v}}$ in place of the effective radius $R_{\mathrm{eff}}$.o This is because our calculated $R_{\mathrm{eff}}$ is not necessarily equivalent to the measure of effective radius typically used for HSW profiles in other works, most likely due to our comparably low physical resolution or our exclusion of over-dense cells.
@@
In addition to the statistics usually obtained from $N$-body simulations or observation, we studied a set of general-relativistic void statistics. In particular, the fluid-intrinsic expansion $\Theta$, the fluid-intrinsic curvature scalar $\tensor\tensor[]{\mathcal{R}}{}$, and the kinetic curvature parameter $\Omega_{\mathrm{K}}$. We found that, as expected, these quantities vary within the void. Numerical relativity simulations such as those we use here allow us to study the general-relativistic qualities of voids without imposing symmetries on the metric tensor. The expansion rate is, on average, 27% greater than the large-scale average of $\Theta=3\bar{H}$ at the centre of voids in the 4 $h^{-1}$ Mpc resolution simulation (15% in the 12 $h^{-1}$ Mpc resolution simulation), and reaches a minimum average value of 97% (91%) of the large-scale average in the 4 $h^{-1}$ Mpc (12 $h^{-1}$ Mpc) resolution simulation.
We find that voids are negatively curved, with the kinetic curvature parameter $\Omega_{\mathrm{K}}$ being 0.8 and 0.6 at the centre of voids at 4 $h^{-1}$ Mpc and 12 $h^{-1}$ Mpc resolution, respectively, while the walls are positively curved, with $\Omega_{\mathrm{K}}$ of -0.55 and -0.19, respectively. We would reasonably expect these values to get even larger in magnitude with a higher (and thus more realistic) physical resolution. Our results thus represent a lower limit on the expansion and curvature of voids in a matter-dominated model universe. This is due to the limitation of our simulations using a continuous fluid approximation; which restricts the smallest physical scales we can sample.
These results are consistent with Timescape model estimates based solely on the angular scales of the CMB acoustic peaks in the 2013 Planck data release, namely $\bar{\Omega}_{\mathrm{K}}=0.862^{+0.024}_{-0.032}$, $\bar{\Omega}_{\mathrm{M}}=0.167^{+0.036}_{-0.037}$, $\bar{\Omega}_{\mathcal{Q}}=-0.0293^{+0.0033}_{-0.0036}$ at present, $z=0$ (see Table 1 and Fig. 1 of Duley et al., 2013). These develop from an initial particle horizon volume average void fraction $f_{{\rm v}i}\mathop{\sim}\limits 2\times 10^{-5}$ at the last scattering surface, when $\bar{\Omega}_{Ki}\mathop{\sim}\limits 7\times 10^{-5}$ and $\bar{\Omega}_{{\cal Q}i}\mathop{\sim}\limits-1\times 10^{-5}$.
@@
While NR simulations in cosmology are not new, our results, quantitatively consistent with backreaction, required the development of the correct analysis tools. The signature of nonzero spatial curvature and its impact on $\Omega_{\mathrm{K}}$ in Figs. 8 and 9 is striking. Its consistency with the kinetic curvature interpretation of cosmic voids raises important questions about distinguishing peculiar velocities from the background, which are central to the backreaction debate. These questions raise the possibility that large angle anomalies observed in the CMB and whole sky catalogues of distant sources (for a recent review, see Aluri et al., 2023) might be resolved within general relativity. We leave delving into these questions for future work.
We emphasise that Newtonian simulations enforce an FLRW expansion of space-time throughout the simulation domain. Such simulations are therefore not well suited to study the relativistic aspects of space-time inside voids.
@@
In this work we have explicitly demonstrated the spatial fluctuations of expansion and curvature in void regions in general relativity. However, the impact of these variations remains to be thoroughly explored. While we find qualitatively similar density profiles to Newtonian results, a strict comparison would require a treatment of matter as collisionless particles alongside NR (or a comparison with a similar fluid-based Newtonian code). The ET has recently been combined with smoothed particle hydrodynamics (SPH; Magnall et al., 2023), however, this adaptation is still being tested for simulations with realistic initial data. In future work, we aim to perform similar simulations using this setup and explore a robust comparison of void statistics between Newtonian gravity and GR.
@@
In addition to our approximation of a continuous fluid, the most important caveats to our results are as follows. We use data from a single spatial slice of each simulation to find the voids. While this is common for void-finding in simulated data it limits our ability to compare with observation, as voids are observed on our past light cone. General-relativistic ray tracing in simulations like those presented here (Macpherson, 2023) would enable us to find the voids in redshift space, allowing for a more direct comparison to observation.
Secondly, our simulations omit a cosmological constant in the evolution. This influences the growth of structures in the model universe, typically resulting in larger density contrasts than a comparable model with $\Lambda$. Such an omission will impact the void statistics in the simulation. We are currently working on implementing this into the Einstein Toolkit, which will allow for closer comparison to Newtonian $N$-body simulations.
@@ Acknowledgements
We thank Nico Hamaus for helpful discussion and comments, and members of the UC Gravity group for their feedback and suggestions.
Support for HJM was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51514.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555.
The simulations used in this work were performed on the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.
The authors wish to acknowledge the use of the University of Canterbury’s Gravity cluster, an inter-faculty HPC resource shared between the School of Mathematics and Statistics and the School of Physical and Chemical Sciences, for the post-processing of simulation data.
@@ Appendix A Resolution study
Fig. 13 shows the stacked mean radial density contrast as given in Fig. 3 (dashed curve) along with the equivalent for the $128^{3}$ simulation (solid curve). Fig. 14 does the same for the three other scalars, $\Theta$, $\tensor\tensor[]{\mathcal{R}}{}$, and $\Omega_{\mathrm{K}}$, from Figs. 7, 8, and 9 respectively. The closeness of the two curves in each of these figures shows that numerical error is not contributing significantly.

### v2

@@ 4.5 Boundary distance profiles
We also produce profiles using this approach, using the set of centres of wall cells adjacent to the interior cells of a given void as the boundary for that void, and define distance from the boundary as the distance to the closest point in that set. We then produce an average of the profiles of 2000 randomly-selected voids (rather than all of them, for computational reasons).
@@
Figs. 10 and 11 show the averaged profile of $\delta$ top panels), $\Theta$ (middle panels), and $\tensor\tensor[]{\mathcal{R}}{}$ (bottom panels) in the two simulations as a function of distance from the void boundary. The extremum that appears in the walls for each variable has greater magnitude than for the corresponding stacked spherical profile. Likewise, the gradient near the walls is greater. These properties are to be expected, as the walls are all aligned at distance $d=0$, rather than scattered at varying radial distances around $r=R_{\mathrm{eff}}$.
@@ 5 Conclusions
In this work, we explored void statistics in non-linear general relativity for the first time. We have implemented the Watershed Einstein Topography (WET) void finder for use with simulation data output by the Einstein Toolkit. Our simulations approximate matter as a continuous fluid, which necessitates differences in the implementation of this void finder compared to publicly-available void finders.
Having identified the voids in two simulations of different physical resolution, we studied the void size function, void fraction, and stacked radial density profiles. We fit HSW profiles (Hamaus et al., 2014) to our stacked radial density profiles for each simulation. As discussed in Section 4.4, we alter this process by fitting a parameter $R_{\mathrm{v}}$ in place of the effective radius $R_{\mathrm{eff}}$. This is because our calculated $R_{\mathrm{eff}}$ is not necessarily equivalent to the measure of effective radius typically used for HSW profiles in other works, likely due to our comparably low physical resolution or exclusion of over-dense cells.
@@
In addition to the statistics usually obtained from $N$-body simulations or observation, we studied a set of general-relativistic void statistics. In particular, the fluid-intrinsic expansion $\Theta$, the fluid-intrinsic curvature scalar $\tensor\tensor[]{\mathcal{R}}{}$, and the kinetic curvature parameter $\Omega_{\mathrm{K}}$. We found that, as expected, these quantities vary within the void. Numerical relativity simulations such as those we use here allow us to study the general-relativistic qualities of voids without imposing symmetries on the metric tensor. The expansion rate is, on average, 27% greater than the large-scale average of $\bar{\Theta}=3\bar{H}$ at the centre of voids in the 4 $h^{-1}$ Mpc resolution simulation (15% in the 12 $h^{-1}$ Mpc resolution simulation), and reaches a minimum average value of 97% (91%) of the large-scale average in the 4 $h^{-1}$ Mpc (12 $h^{-1}$ Mpc) resolution simulation. We find that voids are negatively curved, with the kinetic curvature parameter $\Omega_{\mathrm{K}}$ being 0.83 and 0.61 at the centre of voids at 4 $h^{-1}$ Mpc and 12 $h^{-1}$ Mpc resolution, respectively, while the walls are positively curved, with $\Omega_{\mathrm{K}}$ of -0.58 and -0.18, respectively. We would reasonably expect these values to get even larger in magnitude with a higher (and thus more realistic) physical resolution. Our results thus represent a lower limit on the expansion and curvature of voids in a matter-dominated model universe. This is due to the limitation of our simulations using a continuous fluid approximation; which restricts the smallest physical scales we can sample.
While NR simulations in cosmology are not new, being able to quantify small-scale expansion and curvature in physically relevant regions – i.e., cosmic voids – required the development of appropriate void-finding tools. Although $\Omega_{\mathrm{K}}$ is large near void centres, if one takes averages over randomly chosen regions that contain both voids and walls, its magnitude is small (Macpherson et al., 2019). This can be seen at large values of $r/R_{\mathrm{eff}}$ in the radial profiles and large distances from the void boundary in the boundary distance profiles. It is necessary to unambiguously identify void regions, as we have done here, to see this signal clearly in averages on these scales.
@@
The consistency of Fig. 8 with the kinetic curvature interpretation of cosmic voids raises important questions about distinguishing peculiar velocities from the background and is central to many debates in observational cosmology. These questions raise the possibility that large angle anomalies observed in the CMB and whole sky catalogues of distant sources (for a recent review, see Aluri et al., 2023) might be resolved within general relativity. Work is in progress to address this by examining the detailed properties of the structure dipole in our simulations (Williams et al., in prep).
We emphasise that most cosmological simulations enforce an FLRW expansion of space-time throughout the simulation domain, serving as checks of the standard cosmology. Analytic perturbation methods on FLRW backgrounds have identified gauges in which Newtonian $N$-body results can be embedded (Fidler et al., 2016, 2017), as well as consistency with post-Newtonian results in the presence of large density contrasts (Clifton et al., 2020). However, FLRW has at least three distinct notions of homogeneity and isotropy corresponding to the classic foliation and gauge choices of Bardeen (1980) and their generalizations (Bičák et al., 2007). Possible differences from FLRW expectations are therefore not surprising once strict homogeneity and isotropy of the average evolution is relaxed.
@@
In this work we have explicitly demonstrated the spatial fluctuations of expansion and curvature in void regions in GR. However, the impact of these variations remains to be thoroughly explored. While we find qualitatively similar density profiles to Newtonian results, we have not performed a direct comparison to an equivalent Newtonian model universe. A strict comparison would require a treatment of matter as collisionless particles alongside NR (or a comparison with a similar fluid-based Newtonian code) to minimise numerical artefacts in the comparison. The ET has recently been combined with SPH (Magnall et al., 2023), however, this adaptation is still being tested for simulations with realistic initial data.
While some relativistic effects have been calculated from purely Newtonian simulations (e.g., Milillo et al., 2015; Thomas et al., 2015a; Borzyszkowski et al., 2017) – including some which vanish in the Newtonian limit (e.g., Bruni et al., 2014; Thomas et al., 2015b; Tram et al., 2019) – curvature and expansion profiles inside voids have not yet been studied. Further, such variations in curvature are unable to impact the large-scale expansion of Newtonian simulations due to the implementation of the Friedmann equation. However, calculating curvature and expansion profiles in post-processing is in principle straightforward and it would be interesting to explore how well such profiles can be reliably predicted from density profiles alone. Such an analysis could allow for the generation of general-relativistic profiles from Newtonian $N$-body simulations or observational catalogues. However, this would first require a rigorous comparison of our density profiles to a Newtonian analogue.
The magnitude of $\Omega_{\mathrm{K}}$ we find in void centres is of the same order as that found by Duley et al. (2013, table 1, fig. 1) for the timescape model fit to the angular scales of the CMB acoustic peaks in the 2013 Planck data release. Since more than one time parameter is associated with the present epoch in the timescape model, these being related to particular averaging domains, comparing NR simulations with observation require subtle reinterpretation of standard procedures. Consistency with previous analyses is likely to involve a reinterpretation of the time-dependent reparameterization of Adamek et al. (2016, Sec. 5.3) and is left to future work (Wiltshire et al., in prep). Such studies also necessitate careful analysis of the backreaction density, $\Omega_{\mathcal{Q}}$. Previous analysis found that it cancelled well below the numerical error for averages over the whole simulation box (Macpherson et al., 2019). This may be related to the periodic boundary conditions enforcing a torus topology of the spatial hypersurfaces. The impact of this condition on the size of the backreaction effect has not been studied in GR. However, it is well-known to force backreaction to zero on large scales in Newtonian gravity (Buchert & Ehlers, 1997).
@@
In addition to our approximation of a continuous fluid and periodic boundary conditions, the most important caveats to our results are as follows. We use data from a single spatial slice of each simulation to find the voids. While this is common for void-finding in simulated data, it limits our ability to compare with observation, as voids are observed on our past light cone. General-relativistic ray tracing in simulations like those presented here (Macpherson, 2023) would enable us to find the voids in redshift space, allowing for a more direct comparison to observation. Secondly, our simulations omit a cosmological constant in the evolution. This influences the growth of structures in the model universe, typically resulting in larger density contrasts than a comparable model with $\Lambda$. Such an omission will impact the void statistics in the simulation. We are working on implementing this into the Einstein Toolkit, which will allow for closer comparison to Newtonian $N$-body simulations.
@@ Acknowledgements
We thank the anonymous referee, whose careful reading improved the clarity of our manuscript. We also thank Nico Hamaus for helpful discussion and comments, and members of the UC Gravity and Cosmology group for their feedback and suggestions. Support for HJM was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51514.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. DLW and CS are supported by the Marsden Fund administered by the Royal Society of New Zealand, Te Apārangi under grants M1271 and M1278. The simulations used in this work were performed on the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure. We acknowledge the use of the University of Canterbury’s Gravity cluster, an inter-faculty HPC resource shared between the School of Mathematics and Statistics and the School of Physical and Chemical Sciences, for the post-processing of simulation data.
@@ Appendix A Resolution study
Fig. 12 shows the stacked mean radial density contrast for 2000 randomly-selected voids in the $256^{3}$ cell, 4 $h^{-1}$ Mpc resolution simulation (dashed curve) along with the equivalent for the $128^{3}$ simulation (solid curve). Fig. 13 does the same for the three other scalars, $\Theta$, $\tensor\tensor[]{\mathcal{R}}{}$, and $\Omega_{\mathrm{K}}$, from Figs. 6, 7, and 8 respectively. The closeness of the two curves in each of these figures shows that numerical error is not contributing significantly.